OpenStudy (anonymous):

Let θ be an angle in quadrant 2 such that cos(θ)=-5/13. Find the exact values of cscθ and tanθ.

OpenStudy (perl):



OpenStudy (anonymous):

Well find the Opposite, Adjacent and Hypotenuse with the pythagorean theorem

OpenStudy (anonymous):

which is one way

OpenStudy (anonymous):

you already know the Hypotenuse and the Adjacent since cos = A/H so, a = -5 and H = 13

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\]

OpenStudy (anonymous):

\[c^2 - a^2 = b^2\]

OpenStudy (anonymous):

so.. \[13^2 - 5^2 = b^2\]

OpenStudy (anonymous):

(b is the opposite side)

OpenStudy (anonymous):

\[169 - 25 = 144\]

OpenStudy (anonymous):

\[\sqrt{144} = 12\]

OpenStudy (anonymous):

now you know that the opposite side = 12

OpenStudy (anonymous):

so A = -5, H = 13, O = 12

OpenStudy (anonymous):

tan = O/A and csc = H/O